The Depth of penetration of EM wave in medium having conductivity σ at a frequency of 1 MHz is 25 cm. The depth of penetration at a frequency of 4 MHz will be


1
Expert's answer
2021-01-12T12:15:12-0500

Let's write the formula for the depth of penetration at frequency "f_1":


"\delta_1=\sqrt{\dfrac{1}{\pi\sigma f_1\mu}},"

here, "\delta_1=25\ cm" is the depth of penetration at frequency "f_1=10^6\ Hz", "\sigma" is the conductivity of the material, "\mu" is the permeability.

Similarly, we can write the formula for the depth of penetration at frequency "f_2 = 4\cdot10^6\ Hz":


"\delta_2=\sqrt{\dfrac{1}{\pi\sigma f_2\mu}}."

Let's divide "\delta_1"by "\delta_2", we have:


"\dfrac{\delta_1}{\delta_2}=\sqrt{\dfrac{f_2}{f_1}}."

From this formula, we can find "\delta_2":


"\delta_2=\delta_1\sqrt{\dfrac{f_1}{f_2}}=25\ cm\cdot\sqrt{\dfrac{10^6\ Hz}{4\cdot10^6\ Hz}}=12.5\ cm."

Answer:

"\delta_2=12.5\ cm."